4) From A, B lies 11 km away on a bearing of 0410 and C lies 8 km away on a bearing of 3410. Find

a) The distance between B and C.

b) The bearing of B from C.
1
Expert's answer
2020-02-06T09:18:25-0500


From the figure you can see that the angle between C and B (green angle) is equal to

(360"\degree" -341"\degree" ) + 41"\degree" = 60"\degree"




We get the triangle, let call side AB - c, AC-b, BC-a


1)

So,using the law of cosines:

"a^2 = b^2 + c^2 - 2 bc\times cos A\\
a^2 = 8^2 + 11^2 - 2 (8)(11)\times cos 60\degree\\
a^2 = 97,\\ a = 9.8 km"

BC=9.8 km


2)

From the law of Sines find the angle C at point C

"\frac{9.8}{sin60 \degree} = \frac{11}{sin C}\\ sin C = \frac{sin60 \degree \times 11}{9.8}=0.972\\

 

C=sin^{-1}(0.972 )= 75.3 \degree"


The bearing of B from C is the angle formed by the line joining C and B and rotating about C. By Geometry this angle is

180° - (C + 19°) = 180°-(75.3°+19°)=85.7 "\degree"